# Even Faster Algorithm for the Chamfer Distance

### v1

@@ 5 Rejection Sampling
Let $X:=\{x_{i}\}_{i\in[t]}$ be a set of $t$ samples drawn from a $f$-chamfer distribution $\mathcal{P}$. Fix $h=h(n,d,\varepsilon)\geq 1$. Given an arbitrarily $\tilde{\mathsf{opt}}_{x_{i}}$ for every $x_{i}$ that satisfies $\mathsf{opt}_{x_{i}}\leq\tilde{\mathsf{opt}}_{x_{i}}\leq h\cdot\mathsf{opt}_{x_{i}}$, then for any $10<\kappa<$ 1,
@@
[…]
frametitle= Chamfer-Estimate , skipabove=, skipbelow=, frametitleaboveskip=-0.7frametitlealignment= {mdframed}
Input: Two subsets $A,B$ of a metric space $(\mathbb{R}^{d},\|\cdot\|_{1})$ of size $n$, a parameter $\varepsilon>0$, and a parameter $q\in\mathbb{N}$.
Output: An estimated value $\tilde{\mathsf{CH}}(A,B)\in\mathbb{R}$.
  1. 1.
    Execute the algorithm $\texttt{QuadTree}(A,B)$, and let the output be a set of values $\{\mathcal{D}_{a}\}_{a\in A}$ which always satisfy $\mathcal{D}_{a}\geq\mathsf{opt}_{a}$. Let $\mathcal{D}:=\sum_{a\in A}\mathcal{D}_{a}$.
  2. 2.
    Construct a probability distribution $\mathcal{P}$ supported on $A$ such that for every $a\in A$, $\mathcal{P}(a)=\frac{\mathcal{D}_{a}}{\mathcal{D}}$. For $i\in[q]$, sample $x_{i}\sim\mathcal{P}$.
  3. 3.
    Execute the algorithm $\texttt{Tournament}(\{x_{i}\}_{i\in[q]},B)$, and let the output be a set of values $\{\mathcal{D}^{\prime}_{x_{i}}\}_{i\in[q]}$ which always satisfy $\mathcal{D}^{\prime}_{x_{i}}\geq\mathsf{opt}_{x_{i}}$. Let $\mathcal{D}^{\prime}:=\sum_{i\in[q]}\frac{\mathcal{D}^{\prime}_{x_{i}}}{\mathcal{P}(x_{i})}/q$ and denote $\mathcal{P}^{\prime}(a):=\frac{\mathcal{D}^{\prime}_{a}}{\mathcal{D}^{\prime}}$ (which is well-defined only if $a=x_{i}$ for some $i\in[q]$).
  4. 4.
    Define
    $M:=\max_{i\in[q]}\frac{\mathcal{P}^{\prime}(x_{i})}{\mathcal{P}(x_{i})}.$
    For each $i\in[q]$, mark $x_{i}$ as accepted with probability $\frac{\mathcal{P}^{\prime}(x_{i})}{M\cdot\mathcal{P}(x_{i})}$.
    If the number of accepted $x_{i}$ is less than $s=10/\varepsilon^{2}$ then output Fail and exit the algorithm. Otherwise, collect the first $s$ accepted $x_{i}$ as a set $S:=\{s_{j}\}_{j\in[s]}$.
  5. 5.
    Compute $\mathsf{opt}_{s_{j}}$ for each $j$. Output
    $\tilde{\mathsf{CH}}(A,B):=\sum_{j\in[s]}\frac{\mathsf{opt}_{s_{j}}}{\mathcal{P}^{\prime}(s_{j})}/s.$
@@
Under Condition 5.2, with probability at least $9/10$, $\mathcal{P}$ is a $(1/50{{D}})$-Chamfer Distribution.
@@
Analysis of $S$: We now show that the set $S$ on Line 4 collects enough samples (thus the algorithm does not fail) and is equivalent to sampling from a $\mathcal{O}(1)$-Chamfer distribution $\mathcal{Q}$. We note that the algorithm, in fact, only knows a $(1/50{{D}})$-Chamfer distribution $\mathcal{P}$ and probabilities $\mathcal{P}^{\prime}(x_{i})$ for $\{x_{i}\}_{i\in[q]}$, so it cannot explicitly sample from such $\mathcal{Q}$. Nevertheless, by a standard analysis of rejection sampling, we show that $S$ “simulates” sampling from $\mathcal{Q}$.
@@
This is dominated by the runtime of QuadTree, Tournament, and the time of computing $\mathsf{opt}_{s_{j}}$ on Line 5. $\texttt{QuadTree}(A,B)$ runs in $\mathcal{O}(nd\log\log n)$ time and Tournament ($(\{x_{i}\}_{i\in[q]},B)$ runs in $\mathcal{O}(nd(\log\log n+\log\frac{1}{\varepsilon})/\varepsilon^{2})$ time. Finally, the brute-force search for $\mathsf{opt}_{s_{j}}$ for $j\in[10/\varepsilon^{2}]$ takes $\mathcal{O}(nd/\varepsilon^{2})$ time. ∎
@@ Acknowledgements
Ying Feng was supported by an MIT Akamai Presidential Fellowship. Piotr Indyk was supported in part by the NSF TRIPODS program (award DMS-2022448).
@@
[…]
@@
[…]
@@
[…]
$\leq\frac{1}{2}-c\cdot f_{U}(3/2)$ for $20<c<1/$ 2
@@
Similarly, we can get $\text{\bf Pr}\bigl{[}|v_{i}\cdot(x-y)|<(1-c)\lVert x-y\rVert_{1}\bigr{]}<\frac{1}{2}-c/10$. For $i\in[r]$, let $\mathcal{I}_{i}$ be an indicator variable that equals $1$ if $|v_{i}\cdot(x-y)|<(1-c)\lVert x-y\rVert_{1}$ and equals 00 otherwise. By Hoeffding’s bound,

### v2

@@ 5 Rejection Sampling
Let $X:=\{x_{i}\}_{i\in[t]}$ be a set of $t$ samples drawn from a $f$-chamfer distribution $\mathcal{P}$. Fix $h=h(n,d,\varepsilon)\geq 1$. Given an arbitrarily $\tilde{\mathsf{opt}}_{x_{i}}$ for every $x_{i}$ that satisfies $\mathsf{opt}_{x_{i}}\leq\tilde{\mathsf{opt}}_{x_{i}}\leq h\cdot\mathsf{opt}_{x_{i}}$, then for any $0<\kappa<1$,
@@
Input: Two subsets $A,B$ of a metric space $(\mathbb{R}^{d},\|\cdot\|_{1})$ of size $n$, a parameter $\varepsilon>0$, and a parameter $q\in\mathbb{N}$. Output: An estimated value $\tilde{\mathsf{CH}}(A,B)\in\mathbb{R}$. 1. Execute the algorithm $\texttt{QuadTree}(A,B)$, and let the output be a set of values $\{\mathcal{D}_{a}\}_{a\in A}$ which always satisfy $\mathcal{D}_{a}\geq\mathsf{opt}_{a}$. Let $\mathcal{D}:=\sum_{a\in A}\mathcal{D}_{a}$. 2. Construct a probability distribution $\mathcal{P}$ supported on AA such that for every $a\in A$, $\mathcal{P}(a)=\frac{\mathcal{D}_{a}}{\mathcal{D}}$. For $i\in[q]$, sample $x_{i}\sim\mathcal{P}$. 3. Execute the algorithm $\texttt{Tournament}(\{x_{i}\}_{i\in[q]},B)$, and let the output be a set of values $\{\mathcal{D}^{\prime}_{x_{i}}\}_{i\in[q]}$ which always satisfy $\mathcal{D}^{\prime}_{x_{i}}\geq\mathsf{opt}_{x_{i}}$. Let $\mathcal{D}^{\prime}:=\sum_{i\in[q]}\frac{\mathcal{D}^{\prime}_{x_{i}}}{\mathcal{P}(x_{i})}/q$ and denote $\mathcal{P}^{\prime}(a):=\frac{\mathcal{D}^{\prime}_{a}}{\mathcal{D}^{\prime}}$ (which is well-defined only if $a=x_{i}$ for some $i\in[q]$). 4. Define $M:=\max_{i\in[q]}\frac{\mathcal{P}^{\prime}(x_{i})}{\mathcal{P}(x_{i})}.$ For each $i\in[q]$, mark $x_{i}$ as accepted with probability $\frac{\mathcal{P}^{\prime}(x_{i})}{M\cdot\mathcal{P}(x_{i})}$. If the number of accepted $x_{i}$ is less than $s=10/\varepsilon^{2}$ then output Fail and exit the algorithm. Otherwise, collect the first $s$ accepted $x_{i}$ as a set $S:=\{s_{j}\}_{j\in[s]}$. 5. Compute $\mathsf{opt}_{s_{j}}$ for each $j$. Output $\tilde{\mathsf{CH}}(A,B):=\sum_{j\in[s]}\frac{\mathsf{opt}_{s_{j}}}{\mathcal{P}^{\prime}(s_{j})}/s.$
@@
Under Condition 5.2, with probability at least $9/10$, $\mathcal{P}$ is a (1/50D)(1/50{{D}})-Chamfer Distribution.
@@
Analysis of $S$: We now show that the set $S$ on Line 4 collects enough samples (thus the algorithm does not fail) and is equivalent to sampling from a $\mathcal{O}(1)$-Chamfer distribution $\mathcal{Q}$. We note that the algorithm, in fact, only knows a (1/50D)(1/50{{D}})-Chamfer distribution $\mathcal{P}$ and probabilities $\mathcal{P}^{\prime}(x_{i})$ for $\{x_{i}\}_{i\in[q]}$, so it cannot explicitly sample from such $\mathcal{Q}$. Nevertheless, by a standard analysis of rejection sampling, we show that $S$ “simulates” sampling from $\mathcal{Q}$.
@@
This is dominated by the runtime of QuadTree, Tournament, and the time of computing $\mathsf{opt}_{s_{j}}$ on Line 5. $\texttt{QuadTree}(A,B)$ runs in $\mathcal{O}(nd\log\log n)$ time and Tournament $(\{x_{i}\}_{i\in[q]},B)$ runs in $\mathcal{O}(nd(\log\log n+\log\frac{1}{\varepsilon})/\varepsilon^{2})$ time. Finally, the brute-force search for $\mathsf{opt}_{s_{j}}$ for $j\in[10/\varepsilon^{2}]$ takes $\mathcal{O}(nd/\varepsilon^{2})$ time. ∎
@@ Acknowledgements
Ying Feng was supported by an MIT Akamai Presidential Fellowship. Piotr Indyk was supported in part by the NSF TRIPODS program (award DMS-2022448).The authors would like to thank Anders Aamand for helpful comments that helped simplify the algorithm in Section 4.
@@
In our algorithm, we assumed that all points in input sets $A,B$ are integers in $\{1,2,\cdots,$ ${\rm poly}(n)\}^{d}$. Here, we show that this is without loss of generality, as long as all coordinates of the original input are $w$-bit integers for arbitrary $w\geq\log n$ in a unit-cost RAM with a word length of $w$ bits.
Section $A.3$ of [BIJ+23] gives an efficient reduction from real inputs to the case that
@@
Similar to (the proof of Lemma $A.3$ in) [BIJ+23], we sample a vector $v\sim\mathsf{Cauchy}(0,1)$, which can be discretized to $\mathcal{O}(\log n)$-bit precision following [Ind06]. We then compute the inner products $\{v\cdot a\}_{a\in A}$ and $\{v\cdot b\}_{b\in B}$. The distribution of $v\cdot a-v\cdot b$ follows $\mathsf{Cauchy}(0,\lVert a-b\rVert_{1})$ by the $1$-stability property of Cauchy’s. So we have that for every $a\in A$ and $b\in B$,
@@
$\text{\bf Pr}\bigl{[}U>1+c\bigr{]}$ $=\frac{1}{2}-\int^{1+c}_{1}f_{U}(u)du$
$\leq\frac{1}{2}-c\cdot f_{U}(3/2)$ for $0<c<1/2$
@@
Similarly, we can get $\text{\bf Pr}\bigl{[}|v_{i}\cdot(x-y)|<(1-c)\lVert x-y\rVert_{1}\bigr{]}<\frac{1}{2}-c/10$. For $i\in[r]$, let $\mathcal{I}_{i}$ be an indicator variable that equals $1$ if $|v_{i}\cdot(x-y)|<(1-c)\lVert x-y\rVert_{1}$ and equals 0 otherwise. By Hoeffding’s bound,